OpenStudy (anonymous):

how many real solutions does the equation f(x)=x^2+x+8 have?

OpenStudy (anonymous):

wont have any by either of two methods a) try to solve by completing the square and see that you end up with a negative number on the right hand side or b) compute \(b^2-4ac=1^2-4\times 1\times 8\) and see that the answer is negative, so no solution

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw